# LRQ: Optimizing Post-Training Quantization for Large Language Models by Learning Low-Rank Weight-Scaling Matrices

### v1

@@ Abstract
With the commercialization of large language models (LLMs), weight-activation quantization has emerged to compress and accelerate LLMs, achieving high throughput while reducing inference costs. However, existing post-training quantization (PTQ) techniques for quantizing weights and activations of LLMs still suffer from non-negligible accuracy drops, especially on massive multitask language understanding. To address this issue, we propose Low-Rank Quantization (LRQ) a simple yet effective post-training weight quantization method for LLMs that reconstructs the outputs of an intermediate Transformer block by leveraging low-rank weight-scaling matrices, replacing the conventional full weight-scaling matrices that entail as many learnable scales as their associated weights. Thanks to parameter sharing via low-rank structure, LRQ only needs to learn significantly fewer parameters while enabling the individual scaling of weights, thus boosting the generalization capability of quantized LLMs. We show the superiority of LRQ over prior LLM PTQ works under (i) -bit weight and per-tensor activation quantization, (ii) -bit weight and -bit per-token activation quantization, and (iii) low-bit weight-only quantization schemes. Our code is available at https://github.com/onliwad101/FlexRound_LRQ to inspire LLM researchers and engineers.
@@ 2.3 Low-Rank Quantization
which we refer to as ‘Low-Rank Quantization (LRQ)’. At first, and are initialized to zeros and random values from a normal distribution respectively, and and are set to zero vectors so that starts from a zero matrix like in Eq. 1. Then, , , , , and are learned to minimize in a block-by-block manner. The ablation study on the effect of and in LRQ is presented in Appendix C.
@@ 2.4 Effect of Low-rank Matrices on Generalization of Quantized LLMs
For a calibration sample that is selected from the C4 dataset, even if both FlexRound and LRQ initially start their learning process from the same RTN baseline, LRQ achieves an almost identical accumulated RMSE to FlexRound, as illustrated in Figure 3(a). This observation underscores that the use of low-rank weight-scaling matrices does not pose any noticeable obstacle to the minimization of block reconstruction error on calibration data. For common sense reasoning and MMLU benchmarks that are unseen during the reconstruction stage, however, accumulated RMSE for LRQ is much smaller than that for FlexRound as well as RTN as described in Figure 3(b). This compelling result implies that harnessing the parameter-efficiency of low-rank matrices can facilitate superior generalization on unseen benchmarks. In light of these findings, the incorporation of low-rank matrices into block-wise reconstruction is indeed a pivotal step in enhancing the generalization capability of quantized LLMs. For visual representation across various samples, three figures are incorporated in Appendix D, each depicting the accumulated RMSE for three distinct samples. In addition, we illustrate the sensitivity of the accumulated RMSE to the number of calibration samples in Appendix E.
@@
3 Experiments
In this section, we first explore the influence of the rank in Eq. 2 and the quantity of calibration samples on the performance of LRQ. Next, to verify the effectiveness of LRQ, we compare LRQ with existing state-of-the-art post-training quantization (PTQ) methods for open-source large language models (LLMs) such as Llama and Llama by adopting per-channel asymmetric weight quantization, per-tensor asymmetric static activation quantization, and per-token asymmetric KV cache quantization. For the Llama models in Table 4, however, the accuracy gap on the massive multitask language understanding (MMLU) benchmark between quantized LLMs and their FP16 baselines is observed. In this sense, we also perform experiments on Llama by switching from per-tensor asymmetric static activation quantization to per-token asymmetric activation quantization.
@@ 3 Experiments
We use just a single NVIDIA A100-80GB GPU to quantize all the Llama and Llama models via LRQ. We randomly choose calibration samples with a token length of from the training set of C4 (Raffel et al., 2020), one of the pre-training datasets for Llama. Unless otherwise mentioned, LRQ is applied to all linear layers in both attention and feed-forward modules, and the rank in Eq. 2 is set to for large language models beyond B parameters or to for smaller models so as to reduce the number of learnable parameters by approximately half compared to FlexRound. The exact ratio of the number of learnable parameters in LRQ to the number of pre-trained weights for an intermediate Transformer block of each Llama model is given in Appendix F. All quantized models are evaluated on MMLU (Hendrycks et al., 2021) in the five-shot setting as well as six common sense reasoning benchmarks: BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC easy and challenge (Clark et al., 2018), and OpenBookQA (Mihaylov et al., 2018) in the zero-shot setting. More details are deferred to Appendix H.
@@ 3.1 Ablation Study
3.2 Per-tensor Asymmetric Static Activation Quantization
As meticulously studied in Xiao et al. (2022), per-tensor static activation quantization is hardware-efficient and can be implemented on off-the-shelf GPUs with FasterTransformer, the state-of-the-art Transformer inference framework provided from NVIDIA, to achieve up to inference speed-up and almost halving the memory footprint compared to FP16 baselines. Accordingly, we employ per-tensor asymmetric static activation quantization as well as per-channel asymmetric weight quantization. Moreover, we also quantize the KV cache to -bit with a per-token asymmetric quantization scheme. It is worth noting that for large batch sizes, the KV cache can consume a much larger amount of memory than the model size, thus causing a bottleneck in high-throughput LLM inference. Fortunately, the performance discrepancy before and after per-token asymmetric KV cache quantization is almost insignificant no matter which quantization method is selected, as presented in Appendix G. For this reason, we also additionally utilize per-token asymmetric KV cache quantization. Further experimental details are provided in Appendix H.
Table 1, 2, 3, and 4 reveal the effectiveness of LRQ compared with the state-of-the-art LLM PTQ techniques on common sense reasoning tasks and MMLU, respectively. For common sense reasoning bechmarks, the zero-shot performance of LRQ is almost close to that of FP16 baselines, being superior to that of both SmoothQuant and FlexRound for most of the Llama and Llama models. Not only that, LRQ also considerably outperforms SmoothQuant and FlexRound on MMLU.
3.3 Per-token Asymmetric Activation Quantization
Although LRQ shows better performance than SmoothQuant and FlexRound on both common sense reasoning tasks and MMLU when employing per-tensor asymmetric static activation quantization, there is still the five-shot accuracy gap on MMLU between LRQ and FP16 baselines for Llama as in Table 4. Thus, we also conduct experiments on Llama with a per-token asymmetric activation quantization scheme. More details about experimental settings are given in Appendix H.
@@ 3.3 Per-token Asymmetric Activation Quantization
3.4 Low-bit Weight-only Quantization
As LRQ is designed as a post-training weight quantization technique for LLMs, we also run experiments on weight-only quantization for Llama by comparing LRQ with FlexRound and OmniQuant (Shao et al., 2024), one of the state-of-the-art post-training quantization methods, on common sense reasoning tasks and WikiText2 (Merity et al., 2016). In Table 7, we use per-channel weight quantization instead of group-wise weight quantization to show that LRQ can perform comparably to FP16 baselines on common sense reasoning tasks with -bit per-channel weight quantization. In Table 7, the perplexity of LRQ is almost identical to that of FlexRound and OmniQuant, but the zero-shot performance of LRQ is higher than that of FlexRound and OmniQuant on common sense reasoning tasks. It can be therefore concluded that LRQ is also well-suited for weight-only quantization.
@@ 4 Conclusion
We propose a simple yet effective post-training weight quantization approach for LLMs, LRQ that learns low-rank weight-scaling matrices for block-by-block reconstructing the outputs of an intermediate Transformer block. Using such low-rank matrices, we can decrease the number of learnable parameters effectively while allowing for scaling weights individually due to the sharing of learnable parameters through a low-rank structure, thereby enhancing the generalization of quantized LLMs. We show the superiority of LRQ over prior state-of-the-art LLM PTQ techniques under both weight-activation and weight-only quantization schemes.
@@
To push the limit of post-training weight-activation quantization, two research directions emerge: (i) -bit weight and -bit activation quantization, and (ii) INT4 weight-activation quantization. As explained in Appendix A, Lee et al. (2023a) attempted to quantize LLMs with -bit weight and -bit activation quantization, whereas Wei et al. (2023) and Shao et al. (2024) strived to quantize LLMs with INT6 and even INT4 weight-activation quantization. In this paper, we only deal with the former quantization scheme, -bit weight and -bit activation quantization, which is a limitation of our work.
Like Wei et al. (2023) and Shao et al. (2024), we could also focus on INT4 weight-activation quantization rather than -bit weight and -bit activation quantization in Section 3.3. However, Liu et al. (2023a), an earlier LLM quantization work than Wei et al. (2023) and Shao et al. (2024), already exhibited the non-marginal accuracy degradation of -bit weight and -bit activation quantization despite the fact that Liu et al. (2023a) exploited quantization-aware training, not post-training quantization. For this reason, we prioritize -bit weight and -bit activation quantization over INT4 weight-activation quantization in this paper, thus proposing a new post-training weight quantization technique termed LRQ.
Although we demonstrate the efficacy of LRQ in a -bit weight and -bit activation quantization scheme, the five-shot accuracy gap between LRQ and FP16 baselines is still over percent for Llama B and B on the massive multitask language understanding (MMLU) benchmark. In order to enhance the practicality of weight-activation quantization below INT8 weight-activation quantization, it is essential not only to prevent severe performance degradation associated with INT4 weight-activation quantization but also to narrow such five-shot accuracy gaps to less than percent. We therefore believe that -bit weight and -bit activation quantization is as important as INT4 weight-activation quantization as a future research direction. We also hope that our experimental results would pave the way for the application of -bit weight and -bit activation quantization to LLMs via PTQ in real world.
@@ Appendix A Related Work
Quantization works can be generally categorized into quantization-aware training (QAT) and post-training quantization (PTQ). As QAT can maintain the performance of FP32/FP16 baselines, QAT has been applied to computer vision models (Jung et al., 2019; Esser et al., 2020; Lee et al., 2021). Notwithstanding, there exist many challenges associated with applying QAT to large language models (LLMs) due to the sheer scale of pre-training data and a huge amount of computational resources required for training on the whole pre-training dataset. Although Liu et al. (2023a) presented the possibility of applying QAT to LLMs, unforunately, they did not perform experiments on Llama B, the largest and best performing model among the Llama models, in spite of using a single -GPU node. On the other hand, as Frantar et al. (2023) demonstrated the application of PTQ to LLMs only with a single GPU, many researchers have recently paid attention to PTQ for LLMs.
@@
LLM PTQ can be classified into two categories: LLM weight-only quantization (Frantar et al., 2023; Lin et al., 2023) and LLM weight-activation quantization (Dettmers et al., 2022; Yao et al., 2022; Xiao et al., 2022; Lee et al., 2023b; Wei et al., 2023; Shao et al., 2024). For the former quantization, Frantar et al. (2023) quantized the weights of LLMs into low-bit integers based on layer-wise reconstruction, whereas Lin et al. (2023) did by not counting on reconstruction but per-channel scaling in consideration of both weight and activation magnitudes. Despite the fact that both studies exhibited decent quantization performance, the main benefit of weight-only quantization does not align with serving LLMs with high throughput as delineated in Section 1. In this light, we concentrates on weight-activation quantization.
@@
When it comes to weight-activation quantization, Yao et al. (2022) presented ZeroQuant with a -bit group-wise weight quantization scheme and a -bit per-token activation quantization scheme based on layer-wise knowledge distillation, and Dettmers et al. (2022) proposed LLM.int8() with a -bit per-channel weight quantization scheme and a -bit per-token activation quantization scheme while keeping activation outliers in FP16. As discussed in Xiao et al. (2022), however, ZeroQuant incurs severe accuracy degradation for an open-source LLM, and the inference latency of LLM.int8() can be higher than that of the FP16 baseline. To deal with both issues, Xiao et al. (2022) devised SmoothQuant that can preserve the accuracy of OPT (Zhang et al., 2022) by easing the difficulty of activation quantization and accelerate LLM inference by up to times. Yet, SmoothQuant suffers from non-negligible performance degradation for other open-source models such as Llama and Llama with a -bit per-tensor static activation quantization scheme as illustrated in Figure 1. FlexRound that Lee et al. (2023b) created showed the experimental results of Llama up to B with a -bit per-channel weight quantization scheme and a -bit per-tensor static activation quantization scheme, but FlexRound incurs considerable performance degradation on the massive multitask language understanding (MMLU) benchmark as described in Figure 1(b). Beyond INT8 weight-activation quantization, Lee et al. (2023a) attempted to quantize LLMs with -bit weight and -bit activation quantization, whereas Wei et al. (2023) and Shao et al. (2024) strived to quantize LLMs with INT6 and even INT4 weight-activation quantization.
Appendix B Process of in Eq. 2
Similar to the broadcasting process in Python Numpy, we add , , and .
To be more specific, let be
be
and be
Then, by the broadcasting process, can be expressed as
Appendix C Effect of and in LRQ
@@
Appendix D Figures of Accumulated RMSE on Assorted Samples
Appendix E Sensitivity of Accumulated RMSE to the Number of Calibration Samples

### v2

@@ Abstract
With the commercialization of large language models (LLMs), weight-activation quantization has emerged to compress and accelerate LLMs, achieving high throughput while reducing inference costs. However, existing post-training quantization (PTQ) techniques for quantizing weights and activations of LLMs still suffer from non-negligible accuracy drops, especially on massive multitask language understanding. To address this issue, we propose Low-Rank Quantization (LRQ) a simple yet effective post-training weight quantization method for LLMs that reconstructs the outputs of an intermediate Transformer block by leveraging low-rank weight-scaling matrices, replacing the conventional full weight-scaling matrices that entail as many learnable scales as their associated weights. Thanks to parameter sharing via low-rank structure, LRQ only needs to learn significantly fewer parameters while enabling the individual scaling of weights, thus boosting the generalization capability of quantized LLMs. We show the superiority of LRQ over prior LLM PTQ works under (i) -bit weight and per-tensor activation quantization, (ii) -bit weight and -bit per-token activation quantization, and (iii) low-bit weight-only quantization schemes. Our code is available at Software.
@@ 2.3 Low-Rank Quantization
which we refer to as ‘Low-Rank Quantization (LRQ)’. At first, and are initialized to zeros and random values from a normal distribution respectively, and and are set to zero vectors so that starts from a zero matrix like in Eq. 1. Then, , , , , and are learned to minimize in a block-by-block manner. The ablation study on the effect of and in LRQ is presented in Appendix B.
@@ 2.4 Effect of Low-rank Matrices on Generalization of Quantized LLMs
For a calibration sample that is selected from the C4 dataset, even if both FlexRound and LRQ initially start their learning process from the same RTN baseline, LRQ achieves an almost identical accumulated RMSE to FlexRound, as illustrated in Figure 3(a). This observation underscores that the use of low-rank weight-scaling matrices does not pose any noticeable obstacle to the minimization of block reconstruction error on calibration data. For common sense reasoning and MMLU benchmarks that are unseen during the reconstruction stage, however, accumulated RMSE for LRQ is much smaller than that for FlexRound as well as RTN as described in Figure 3(b). This compelling result implies that harnessing the parameter-efficiency of low-rank matrices can facilitate superior generalization on unseen benchmarks. In light of these findings, the incorporation of low-rank matrices into block-wise reconstruction is indeed a pivotal step in enhancing the generalization capability of quantized LLMs. For visual representation across various samples, three figures are incorporated in Appendix C, each depicting the accumulated RMSE for three distinct samples. In addition, we illustrate the sensitivity of the accumulated RMSE to the number of calibration samples in Appendix D.
@@
3 Experiments
In this section, we first explore the influence of the rank in Eq. 2 and the quantity of calibration samples on the performance of LRQ. Next, to verify the effectiveness of LRQ, we compare LRQ with existing state-of-the-art post-training quantization (PTQ) methods for large language models (LLMs).
@@ 3 Experiments
| Method | # Bits (W/A/KV) | Average | WikiText2 |
| Llama B | |||
| GPTQ | |||
| AWQ | |||
| QuIP | |||
| FlexRound | |||
| LRQ (Ours) |
We use just a single NVIDIA A100-80GB GPU to quantize LLMs via LRQ. We randomly choose calibration samples with a token length of from the training set of C4 (Raffel et al., 2020). Unless otherwise mentioned, LRQ is applied to all linear layers in both attention and feed-forward modules, and the rank in Eq. 2 is set to for large language models beyond B parameters or to for smaller models so as to reduce the number of learnable parameters by approximately half compared to FlexRound. The exact ratio of the number of learnable parameters in LRQ to the number of pre-trained weights for an intermediate Transformer block of each Llama model is given in Appendix J. Quantized models are evaluated on MMLU (Hendrycks et al., 2021) in the five-shot setting or common sense reasoning benchmarks: BoolQ (Clark et al., 2019), PIQA (Bisk et al., 2020), HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), ARC easy and challenge (Clark et al., 2018), and OpenBookQA (Mihaylov et al., 2018) in the zero-shot setting. More details are deferred to Appendix I.
@@ 3.1 Ablation Study
3.2 Per-tensor Static Activation Quantization
As meticulously studied in Xiao et al. (2022), per-tensor static activation quantization is hardware-efficient and can be implemented on off-the-shelf GPUs with FasterTransformer, the state-of-the-art Transformer inference framework provided from NVIDIA, to achieve up to inference speed-up and almost halving the memory footprint compared to FP16 baselines. Accordingly, we employ per-tensor asymmetric static activation quantization as well as per-channel asymmetric weight quantization. Moreover, we also quantize the KV cache to -bit with a per-token asymmetric quantization scheme. It is worth noting that for large batch sizes, the KV cache can consume a much larger amount of memory than the model size, thus causing a bottleneck in high-throughput LLM inference. Fortunately, the performance discrepancy before and after per-token asymmetric KV cache quantization is almost insignificant no matter which quantization method is selected, as presented in Appendix H. For this reason, we also additionally utilize per-token asymmetric KV cache quantization. Further experimental details are provided in Appendix I.
Table 1, 2, 3, and 4 reveal the efficacy of LRQ on common sense reasoning tasks and MMLU. For common sense reasoning tasks, the zero-shot accuracy of LRQ is almost close to that of FP16 baselines, being superior to that of both SmoothQuant and FlexRound for most of the Llama and Llama models. Not only that, LRQ also considerably outperforms SmoothQuant and FlexRound on MMLU.
3.3 Per-token Activation Quantization
Although LRQ shows better performance than SmoothQuant and FlexRound on both common sense reasoning tasks and MMLU when employing per-tensor asymmetric static activation quantization, there is still the five-shot accuracy gap on MMLU between LRQ and FP16 baselines for Llama as in Table 4. Thus, we also conduct experiments on Llama with a per-token asymmetric activation quantization scheme. More details about experimental settings are given in Appendix I.
@@ 3.3 Per-token Activation Quantization
3.4 Per-channel Weight-only Quantization
As LRQ is designed as a post-training weight quantization technique for LLMs, we also run experiments on weight-only quantization for Llama and Llama B (Dubey et al., 2024) on common sense reasoning tasks and WikiText2 (Merity et al., 2016). In Table 7 and 8, we use per-channel weight-only quantization instead of group-wise weight-only quantization. Table 7 and 8 show that the average zero-shot accuracy of LRQ is consistently higher than that of all PTQ methods on common sense reasoning tasks. Specifically, in Table 7, LRQ can perform closely to FP16 baselines on common sense reasoning tasks even with 3-bit per-channel weight-only quantization. Despite the fact that Llama is hard to quantize, Table 8 exhibits that LRQ attains the smallest accuracy drop (less than percent compared to the FP16 baseline) on common sense reasoning tasks. In addition, Figure 5 displays compression and acceleration after quantizing Llama 2 to 4-bit via LRQ, showing that LRQ can perform comparably to FP16 baselines while reducing both latency and model size noticeably.
@@ 4 Conclusion
We propose a simple yet effective post-training weight quantization approach for LLMs, LRQ that learns low-rank weight-scaling matrices for block-by-block reconstructing the outputs of an intermediate Transformer block. LRQ can decrease the number of learnable parameters effectively while allowing for scaling weights individually, thereby enhancing the generalization of quantized LLMs.
@@
To push the limit of post-training weight-activation quantization, two research directions emerge: (i) -bit weight and -bit activation quantization, and (ii) INT4 weight-activation quantization. As explained in Appendix A, Lee et al. (2023a) attempted to quantize LLMs with -bit weight and -bit activation quantization, whereas Wei et al. (2023) and Shao et al. (2024) strived to quantize LLMs with INT6 and even INT4 weight-activation quantization. In this paper, we only deal with the former quantization scheme, -bit weight and -bit activation quantization.
Like Wei et al. (2023) and Shao et al. (2024), we could also focus on INT4 weight-activation quantization rather than -bit weight and -bit activation quantization in Section 3.3. However, Liu et al. (2023a), an earlier LLM quantization work than Wei et al. (2023) and Shao et al. (2024), already exhibited the non-marginal accuracy degradation of -bit weight and -bit activation quantization despite the fact that Liu et al. (2023a) exploited quantization-aware training, not post-training quantization. Furthermore, in terms of serving throughput, Lin et al. (2024) shows the superiority of -bit weight and -bit activation quantization over INT4 weight-activation quantization as well as INT8 weight-activation quantization. For these reasons, we prioritize -bit weight and -bit activation quantization over INT4 weight-activation quantization in this paper.
@@ Appendix A Related Work
Quantization works can be generally categorized into quantization-aware training (QAT) and post-training quantization (PTQ). As QAT can maintain the performance of FP32/FP16 baselines, QAT has been applied to computer vision models (Jung et al., 2019; Esser et al., 2020; Lee et al., 2021). Notwithstanding, there exist many challenges associated with applying QAT to large language models (LLMs) due to the sheer scale of pre-training data and a huge amount of computational resources required for training on the whole pre-training dataset. Although Liu et al. (2023a) presented the possibility of applying QAT to LLMs, unfortunately, they did not perform experiments on Llama B, the largest and best performing model among the Llama models, in spite of using a single -GPU node. Kim et al. (2023) introduced Parameter-Efficient and Quantization-aware Adaptation (PEQA) that fine-tunes only quantization step sizes. While PEQA consumes less computational resources and can perform better than QAT, it still requires a single -GPU node to quantize Llama B.
@@
As Frantar et al. (2023) demonstrated the application of PTQ to LLMs only with a single GPU, many researchers have recently paid attention to PTQ for LLMs. LLM PTQ can be classified into two categories: LLM weight-only quantization (Frantar et al., 2023; Lin et al., 2023; Chee et al., 2023) and LLM weight-activation quantization (Dettmers et al., 2022; Yao et al., 2022; Xiao et al., 2022; Lee et al., 2023b; Wei et al., 2023; Shao et al., 2024). For the former quantization, Frantar et al. (2023); Chee et al. (2023) quantized the weights of LLMs into low-bit integers based on layer-wise reconstruction, whereas Lin et al. (2023) did by not counting on reconstruction but per-channel scaling in consideration of both weight and activation magnitudes. Despite the fact that all these studies exhibited decent quantization performance, the main benefit of weight-only quantization does not align with serving LLMs with high throughput as delineated in Section 1. In this light, we concentrates on weight-activation quantization.
@@
When it comes to weight-activation quantization, Yao et al. (2022) presented ZeroQuant with a -bit group-wise weight quantization scheme and a -bit per-token activation quantization scheme based on layer-wise knowledge distillation, and Dettmers et al. (2022) proposed LLM.int8() with a -bit per-channel weight quantization scheme and a -bit per-token activation quantization scheme while keeping activation outliers in FP16. As discussed in Xiao et al. (2022), however, ZeroQuant incurs severe accuracy degradation for an open-source LLM, and the inference latency of LLM.int8() can be higher than that of the FP16 baseline. To deal with both issues, Xiao et al. (2022) devised SmoothQuant that can preserve the accuracy of OPT (Zhang et al., 2022) by easing the difficulty of activation quantization and accelerate LLM inference by up to times. Yet, SmoothQuant suffers from non-negligible performance degradation for other open-source models such as Llama and Llama with a -bit per-tensor static activation quantization scheme as illustrated in Figure 1. FlexRound that Lee et al. (2023b) created showed the experimental results of Llama up to B with a -bit per-channel weight quantization scheme and a -bit per-tensor static activation quantization scheme, but FlexRound incurs considerable performance degradation on the massive multitask language understanding (MMLU) benchmark as described in Figure 1(b). Beyond INT8 weight-activation quantization, Lee et al. (2023a) attempted to quantize LLMs with -bit weight and -bit activation quantization, whereas Wei et al. (2023) and Shao et al. (2024) strived to quantize LLMs with INT6 and even INT4 weight-activation quantization. However, in terms of serving throughput, Lin et al. (2024) showed the superiority of -bit weight and -bit activation quantization over INT4 weight-activation quantization as well as INT8 weight-activation quantization.
Appendix B Effect of and in LRQ
@@
Appendix C Figures of Accumulated RMSE on Assorted Samples
Appendix D Sensitivity of Accumulated RMSE to the Number of Calibration Samples